OpenStudy (beccab003):

Can someone help me understand how to simplify radicals? I have been through so many problems and I still don't understand. Thanks! Medal to best answer.

OpenStudy (anonymous):

how r u doing so far

OpenStudy (beccab003):

I have been doing practice problems but I have not gotten through one 100% correctly. I get lost in the steps to simplify! I can put a practice problem up if you want so you can help me with the process.

OpenStudy (danjs):

post an example, and i will go through the properties used to simplify

OpenStudy (danjs):

there really are only a couple properties to remember

OpenStudy (beccab003):

\[\sqrt[3]{-128xy ^{4}}\]

OpenStudy (danjs):

remember any root sign also means the quantity underneath raised to the power of 1 over the order of the root... like this..

OpenStudy (danjs):

\[\sqrt[n]{a} = a^{1/n}\]

OpenStudy (danjs):

that is property 1

OpenStudy (beccab003):

Okay, am I supposed to find the perfect square of -128?

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} = (-128xy^4)^{1/3}\]

OpenStudy (beccab003):

oh I see, it has to do with the 3 on the \[\sqrt[3]{?}\]

OpenStudy (beccab003):

the 1/3 that is

OpenStudy (danjs):

yeah that is the cubed root

OpenStudy (danjs):

same as raising the stuff to the (1/3) power

OpenStudy (beccab003):

so, for instance if the number was a 2 like this: \[\sqrt[2]{?}\] it would be 1/2?

OpenStudy (danjs):

The square root is the same as raising to the (1/2) power

OpenStudy (danjs):

lol yep, but for 2, it is just square root, you dont write the 2

OpenStudy (danjs):

only numbers higher than 2

OpenStudy (beccab003):

so, it wouldn't be 1/2?

OpenStudy (danjs):

no i am saying, when it is a 2, you just write \[\sqrt{a} = \sqrt[2]{a}\] usually dont put a 2 in there, only if the number is a higher root than 2 (the square root)

OpenStudy (danjs):

If you see this \[\sqrt{x} = x^{1/2}\] the 2 is implied

OpenStudy (beccab003):

okay. so with the number 4 it would be 1/4?

OpenStudy (danjs):

yeah, the 4th root of something, is the same as something to the 1/4 power

OpenStudy (beccab003):

okay, now back to the example...

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} = (-128xy^4)^{1/3}\]

OpenStudy (danjs):

The thing i usually do , if you want to simplify that, is leave the numbers as roots, and put the variables to the powers, like this....

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} =\sqrt[3]{-128}*(xy^4)^{1/3}\]

OpenStudy (danjs):

good on that part?

OpenStudy (beccab003):

Yes

OpenStudy (danjs):

ok, now for the exponents \[[a^b]^c = a^{b*c}\] a power raised to a power, is the base to the powers multiplied together

OpenStudy (danjs):

that is property 2 , if you want to call it that

OpenStudy (beccab003):

Great! So it's 4*1/3 ??

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} =\sqrt[3]{-128}*(xy^4)^{1/3} = \sqrt[3]{-128}*x^{1 * 1/3}*y^{4 * 1/3}\]

OpenStudy (danjs):

yep, move the power onto both the x and the y, like that above

OpenStudy (beccab003):

4*1/3 = 1.33333

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} =\sqrt[3]{-128}*(xy^4)^{1/3} = \sqrt[3]{-128} * x^{1/3} * y ^{4/3}\]

OpenStudy (danjs):

nah, no decimal approximations, leave it as exact fractions

OpenStudy (beccab003):

OOOH I get it. gotta keep em as fractions.

OpenStudy (beccab003):

What's next boss?

OpenStudy (danjs):

yes, and that is all you can do to the variables, now for the cubed root of -128

OpenStudy (danjs):

\[\sqrt[3]{-128} = \sqrt[3]{-64 * 2}\]

OpenStudy (danjs):

see what i did there, 64 is a perfect cube 4*4*4

OpenStudy (beccab003):

Okay. And I understand that you are supposed to get the number to it's most simple cubed root form. How do you know if it is at it's most simple? Also, how do you know a perfect square is 64?

OpenStudy (danjs):

a perfect cube is 64, 4 ^3 = 4 * 4 * 4

OpenStudy (danjs):

check if the number under the cubed root is a multiple of any of the perfect cubes, 2^3 = 8, or 3^3 = 27, or 4^3 = 64 ...and so on

OpenStudy (danjs):

and you notice 128 is 2 * 4^3

OpenStudy (beccab003):

do you just divide by 2? or did you do 4^3. I run into the most issues when I am trying to find the perfect square.

OpenStudy (danjs):

yeah squares are easier, if it is an even number, just divide by 2 , then 2 again, then you will have 4 * (number) and square root of 4 is 2

OpenStudy (beccab003):

okay that makes it easier.

OpenStudy (danjs):

just write your number out into its factors..like for 128 128 = 2 * 64 = 2 * 2 * 32 = 2 * 2 * 2 * 16 and there you see a perfect cube, 2^3 (3 2's)

OpenStudy (beccab003):

okay

OpenStudy (beccab003):

So we have: \[\sqrt[3]{-64*2}\]

OpenStudy (danjs):

\[\sqrt[3]{-128} = \sqrt[3]{(-4 * 4 * 4 * 2}) = \sqrt[3]{(-4)^3 * 2} = -4*\sqrt[3]{2}\]

OpenStudy (danjs):

you can pull out 7 2's from 128 128 = 2 * 2 * 2 * 2 * 2 * 2 * 2 = 2^7

OpenStudy (danjs):

so you get 2^6 * 2\[\sqrt[3]{128} = \sqrt[3]{2^6 * 2} = 2^{6/3}*\sqrt[3]{2} = 2^2\sqrt[3]{2} = 4 * \sqrt[3]{2}\]

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} = -4 *\sqrt[3]{2}*x^{1/3}*y^{4/3}\]

OpenStudy (danjs):

or you can just leave it as this if you dont want a root sign....

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} = -2^{7/3} *x^{1/3}*y^{4/3}\]

OpenStudy (beccab003):

Okay. Isn't the x and the y pulled over to the left side though?

OpenStudy (danjs):

What do you mean? in front of the numbers?

OpenStudy (danjs):

multiplication can be done in any order

OpenStudy (danjs):

usually you write the constants first, then the variables

OpenStudy (beccab003):

aren't the x and the y variables supposed to be pulled to the left side.

OpenStudy (danjs):

i'm not gettin what you mean

OpenStudy (beccab003):

because this is an example I actually have the right answer to this and it is: \[-4y \sqrt[3]{2xy}\]

OpenStudy (danjs):

i see what they want there ... just take out the perfect cubed variables and leave the rest underneath

OpenStudy (danjs):

\[\sqrt[3]{-128xy^4} = \sqrt[3]{(-2^6 * 2*x * y^3 * y} = -2^2*y* \sqrt[3]{2*x*y}\]

OpenStudy (danjs):

they just took out only the perfect cubes, instead of the whole thing like i did

OpenStudy (beccab003):

alright. can we do one more but this time I'll try to do it to see if I can go through the process?

OpenStudy (danjs):

sure, can you post a new thread, and tag me, and immediately close it out so people dont swarm to it

OpenStudy (beccab003):

yup